# Conformal Prediction with Upper and Lower Bound Models

### v2

@@ 1 Introduction
Optimization technology is instrumental in operating complex systems efficiently, from supply chains, logistics and manufacturing to healthcare and power systems. Nevertheless, it is often intractable to obtain optimal solutions to large-scale, nonlinear and non-convex problems found in industry. This limitation has motivated the use of approximate and heuristic algorithms, which are fast but may return sub-optimal solutions. As a consequence, decision makers face the hurdle of assessing whether a given solution is optimal enough, or whether substantial improvements are possible. This assessment requires computing dual bounds, i.e., mathematical certificates of solution quality, which are typically obtained by solving relaxations boyd2004convex; geoffrion2009lagrangean or by branch-and-bound algorithms wolsey1999integer. However, in real-life applications, these dual bounds may be too loose to be informative, thus leaving practitioners to evaluate solution quality by relying on their domain knowledge and experience, which is prone to errors and may lead to sub-optimal decisions.
To address this challenge, this paper proposes a data-driven approach that evaluates the sub-optimality of a solution by leveraging uncertainty quantification (UQ) and conformal prediction (CP) techniques. Namely, the paper proposes to provide probabilistic guarantees for solution quality in the form of prediction intervals. To the best of the authors’ knowledge, this paper is the first to propose the use of CP to estimate solution sub-optimality in an optimization setting. In doing so, the paper introduces CPUL, a general framework for CP with upper and lower bounds. This contrasts with existing approaches from the optimization literature, which almost exclusively rely on computing valid, but possibly loose, dual bounds. The paper’s setting also contrasts with traditional CP settings that rely on approximate predictors, e.g., mean predictors vovk2005algorithmic or quantile regression models romano2019conformalized which, in the present setting, may not yield valid lower/upper bounds.
@@ 1.1 Problem statement
where $x\in\mathcal{X}\subseteq\mathbb{R}^{p}$ denotes the instance’s parameters, $s\in\mathbb{R}^{n}$ are decision variables, $f_{x}:\mathbb{R}^{m}\rightarrow\mathbb{R}$ is the objective function to be minimized, $g_{x}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ encodes constraints and $\mathcal{S}_{x}=\{g_{x}(s)\leq 0\}$ is the set of feasible solutions, and $\Phi(x)$ is the instance’s optimal value. The paper assumes the availability of a primal-dual pair in the following form. A primal solution $\bar{s}\in\mathcal{S}_{x}$ provides a primal (upper) bound on the optimal value, i.e., $f_{x}(\bar{s})\geq\Phi(x)$. Primal solutions are obtained using either heuristics of exact algorithms. Conversely, a dual (lower) bound $\psi\in\mathbb{R}$ is guaranteed to be smaller than the optimal value, i.e., $\psi\leq\Phi(x)$. Dual bounds are obtained from relaxations and/or branch-and-bound algorithms, and provide a mathematical certificate on the quality of a primal solution $\bar{s}$. It is important to note that all modern global optimization algorithms return such a primal-dual pair. It is also possible to obtain a primal-dual pair using only machine learning (ML) models, by combining primal-feasible ML architectures donti2021dc3; chen2023end with the dual lagrangian learning framework introduced in tanneau2024dual.
@@
These predictors inherently form a 100% prediction interval $[\hat{B}^{l}(X),\hat{B}^{u}(X)]$ for $Y$. Recall that, in a constrained optimization context, upper and lower predictors are obtained from primal and dual information, respectively. The goal of this paper is to derive tighter prediction intervals by targeting a $(1-\alpha)$ marginal coverage for $Y$ (where $\alpha\,{\in}\,[0,1]$), rather than perfect coverage.
Formally, consider a dataset of $N$ independent and identically distributed (i.i.d.) observations $(X_{i}\,{,}\,Y_{i})_{i\in[N]}$, where $X_{i}$ is sampled from an underlying distribution $\mathcal{P}$, and $Y_{i}\,{=}\,f(X_{i}),\forall i\,{\in}[N]$. Following the standard procedure of the split CP framework vovk2005algorithmic, divide the $N$ samples into a training and a calibration set, denoted as $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{cal}},$ with corresponding indices $\mathcal{I}_{\text{train}},\mathcal{I}_{\text{cal}}$. Next, consider a new i.i.d. observation $X_{N+1}$, sampled from $\mathcal{P}$. For $\alpha\,{\in}\,[0,1]$, the goal is to refine the initial prediction interval $\tilde{C}(X_{N+1})=[\hat{B}^{l}(X_{N+1}),\hat{B}^{u}(X_{N+1})]$ into a (smaller) prediction interval $\hat{C}(X_{N+1})=[\hat{L}(X_{N+1}),\hat{U}(X_{N+1})]$ that satisfies the marginal coverage guarantee
@@ 2 Background
UQ is crucial for informed decision making, particularly in real-world applications where prediction sets are preferred over point estimates. CP, first proposed by vovk2005algorithmic, is a widely used distribution-free UQ method, valued for its finite-sample coverage guarantees and computational efficiency. This section presents standard CP techniques, including Split CP, Conformal Quantile Regression, and Nested CP. While these methods can be applied to the problem at hand (see Section 5 and Appendix A.4.1), unlike the proposed framework, they are not designed to explicitly leverage valid lower and upper bound predictors.
@@
vovk2005algorithmic; papadopoulos2002inductive; lei2018distribution is one of the most commonly used CP frameworks. Therein, a mean prediction model $\hat{f}$ is first trained using the training data $\mathcal{D}_{train}$. SCP then produces prediction intervals of the form
@@
where $\hat{s}\,{:}\,\mathbb{R}^{d}{\times}\mathbb{R}\,{\rightarrow}\,\mathbb{R}$ is a so-called conformity score, and $\hat{Q}^{\text{SCP}}_{\frac{\alpha}{2}},\hat{Q}^{\text{SCP}}_{1-\frac{\alpha}{2}}$, denote the empirical $\frac{\alpha}{2}$ and $1{-}\frac{\alpha}{2}$ quantiles of the scores $\{\hat{s}(X_{i},Y_{i})\}_{i\in\mathcal{I}_{\text{cal}}}$. Under the exchangeability assumption, the resulting prediction intervals have valid marginal coverage vovk2005algorithmic, i.e.,
@@
Common choices of conformity scores include the residual score $\hat{s}(x,y)\,{=}\,y{-}\hat{f}(x)$ and the absolute residual score $\hat{s}(x,y)\,{=}\,|y{-}\hat{f}(x)|$; the reader is referred to angelopoulos2021gentle; oliveira2024split for a more exhaustive review of SCP and conformity scores. The SCP methodology can be applied to the paper’s setting by replacing $\hat{f}$ in the above construction with either the lower ($\hat{B}^{l}$) or upper bound ($\hat{B}^{u}$) predictor. However, this fails to jointly utilize information from both bounds.
@@
To alleviate SCP’s lack of local adaptivity, CQR romano2019conformalized combines quantile regression models with a conformalization procedure: First, CQR trains quantile regression models to predict the quantiles $\alpha_{\text{lo}}$ and $\alpha_{\text{hi}}$. Then, conformalize as follows
@@
gupta2022nested introduces a unifying CP framework by nested prediction sets. NCP considers a family of nested prediction intervals $\{\hat{C}_{t}\}_{t\in\mathcal{T}\subseteq\mathbb{R}}$, i.e.,
@@
is computed from the calibration set. This framework encompasses all types of conformity scores, while maintaining the theoretical guarantees of standard CP gupta2022nested. For instance, the CQR construction can be cast in the NCP framework by considering
@@
CP techniques have been use to analyze the precision of sketching algorithms, which are increasingly vital tools for handling massive datasets in modern machine learning broder2004network; goyal2012sketch; cormode2018privacy; zhang2014these. sesia2023conformal construct prediction intervals for the frequency of a queried object based on a sketch charikar2002finding. Therein, a deterministic upper bound $\hat{B}^{u}$ provided by cormode2005improved is combined with a trivial lower bound $\hat{B}^{l}=0$ (reflecting the non-negativity of counts) to construct nested intervals. The construction, referred to as “SFD CP” in this paper, is defined as:
@@
sesia2023conformal also introduces an adaptive version of this construction, designed to better account for heteroscedasticity in residuals. The CPUL framework generalizes this approach by combining information from $\hat{B}^{l},\hat{B}^{u}$ in multiple ways; see Section 3.2.
Another related line of work leverages information from multiple models fitted on the training set. For example, liang2024conformal; yang2024selection focus on selecting the model that produces the most efficient prediction intervals. Extending these ideas, the CPUL framework reinterprets the problem in the context of multiple models, simultaneously utilizing both upper and lower bounds to maximize the utility of the available information.
@@ 3 CP With Upper And Lower Bounds
(b) Illustration of Paradoxical Miscoverage (motivation for OMLT): using a constant offset (±$\pm t$) in the NCP construction results in an empty prediction interval where $\hat{B}^{u}(x)-\hat{B}^{l}(x)$ is small.
@@ 3.1 Exploiting valid bounds in NCP
Recall that $\hat{B}^{l},\hat{B}^{u}$ provide valid lower and upper bounds on the target variable, i.e., $\hat{B}^{l}(X)\,{\leq}\,Y\,{\leq}\,\hat{B}^{u}(X)$ always holds. This suggests a simple procedure for strengthening $\hat{C}$ without loss of coverage (as shown by Proposition 3.1) into
@@
$\hat{C}(x)\,{\cap}\,\big{[}\hat{B}^{l}(x),\hat{B}^{u}(x)\big{]},\forall x\in\mathcal{X}$. Then, $\mathbb{P}\big{(}Y_{N+1}\,{\in}\,\tilde{C}(X_{N+1})\big{)}=\mathbb{P}\big{(}Y_{N+1}\,{\in}\,\hat{C}(X_{N+1})\big{)}=1{-}\alpha.$appendixproof} It suffices to prove that $\mathbb{P}\left(Y\in\tilde{C}(X)\right)=\mathbb{P}\left(Y\in\hat{C}(X)\right)$. First note that
$\hat{C}(X)$ $=\hat{C}(X)\cap\mathbb{R}$
$=\hat{C}(X)\cap\left((-\infty,\hat{B}^{l}(X))\cup[\hat{B}^{l}(X),\hat{B}^{u}(X))]\cup(\hat{B}^{u}(X),+\infty)\right)$
and note that $(-\infty,\hat{B}^{l}(X))$, $[\hat{B}^{l}(X),\hat{B}^{u}(X))]$ and $(\hat{B}^{u}(X),+\infty)$ are disjoint. Also note that
$\mathbb{P}\left(Y\in\hat{C}(X)\cap(-\infty,\hat{B}^{l}(X))\right)=\mathbb{P}\left(Y\in\hat{C}(X)\cap(\hat{B}^{u}(X),+\infty)\right)=0,$
because $\hat{B}^{l}(X)\leq Y\leq\hat{B}^{u}(X)$ by definition of $\hat{B}^{l},\hat{B}^{u}$. It then follows that
$\mathbb{P}\left(Y\in\hat{C}(X)\right)$ $=\mathbb{P}\left(Y\in\hat{C}(X)\cap(-\infty,\hat{B}^{l}(X))\right)+\mathbb{P}\left(Y\in\hat{C}(X)\cap[\hat{B}^{l}(X),\hat{B}^{u}(X))]\right)+\mathbb{P}\left(Y\in\hat{C}(X)\cap(\hat{B}^{u}(X),+\infty)\right)$
$=0+\mathbb{P}\left(Y\in\hat{C}(X)\cap[\hat{B}^{l}(X),\hat{B}^{u}(X))]\right)+0$
$=\mathbb{P}\left(Y\in\tilde{C}(X)\right)$
which concludes the proof. It is important to note that Proposition 3.1 holds irrespective of how the original prediction interval $\hat{C}$ is obtained. Theorem 3.1 shows that, if the calibration step is performed using the NCP framework, then there is no loss of performance whether the strengthening is performed before or after calibration. {theoremrep} Consider a family of nested prediction sets $\{\hat{C}_{t}\}_{t\in\mathcal{T}}$, where $\mathcal{T}\,{\subseteq}\,\mathbb{R}$, and let $\hat{\tau}$ be obtained following the NCP calibration step as per Eq. (11). Next, define the family of nested strengthened intervals $\{\tilde{C}_{t}\}_{t\in\mathcal{T}}$, where $\forall x\in\mathbb{R}^{d},\forall t\in\mathcal{T},\tilde{C}_{t}(x)=\hat{C}_{t}(x)\cap\big{[}\hat{B}^{l}(x),\hat{B}^{u}(x)\big{]},$ and let $\tilde{\tau}$ be obtained from Eq. (11). Then, $\hat{\tau}=\tilde{\tau}$ and $\forall x\in\mathbb{R}^{d},$
@@
{appendixproof}
First note that $\{\tilde{C}_{t}\}_{t\in\mathcal{T}}$ is indeed a family of nested intervals that satisfies the NCP assumptions; this follows from the fact that $\tilde{C}_{t}\subseteq\hat{C}_{t},\forall t$.
Next, using the same argument as the proof of Proposition 3.1,
$\forall i\in\mathcal{I}_{\text{cal}},\mathbf{1}_{\hat{C}_{t}(X_{i})}(Y_{i})=\mathbf{1}_{\tilde{C}_{t}(X_{i})}(Y_{i})$ (16)
Substituting this in Eq. (11) then yields
$\left\{t\,\middle|\,\sum_{i\in\mathcal{I}_{\text{cal}}}\mathbf{1}_{\tilde{C}_{t}(X_{i})}(Y_{i})\geq(1{-}\alpha)(1{+}|\mathcal{I}_{cal}|)\right\}$ =$=\left\{t\,\middle|\,\sum_{i\in\mathcal{I}_{\text{cal}}}\mathbf{1}_{\hat{C}_{t}(X_{i})}(Y_{i})\geq(1{-}\alpha)(1{+}|\mathcal{I}_{cal}|)\right\},$ (17)
which then yields $\hat{\tau}=\tilde{\tau}$ by unicity of the infimum, and $\hat{C}_{\hat{\tau}}(X)\cap[\hat{B}^{l}(X),\hat{B}^{u}(X)]=\tilde{C}_{\tilde{\tau}}(X)$ follows immediately. Theorem 3.1 allows to decouple the calibration step from the strengthening. Thereby, one can perform the calibration step without access to $\hat{B}^{l}$ or $\hat{B}^{u}$, without any loss of coverage. This is particularly valuable when $\hat{B}^{l}$ or $\hat{B}^{u}$ are not available during training/calibration due to, e.g., privacy concerns.

### v3

@@ 1 Introduction
Optimization technology is instrumental in operating complex systems efficiently, from supply chains, logistics and manufacturing to healthcare and power systems. Nevertheless, it is often intractable to obtain optimal solutions to large-scale, nonlinear and non-convex problems found in industry. This limitation has motivated the use of approximate and heuristic algorithms, which are fast but may return sub-optimal solutions. As a consequence, decision makers face the hurdle of assessing whether a given solution is optimal enough, or whether substantial improvements are possible. This assessment requires computing dual bounds, i.e., mathematical certificates of solution quality, which are typically obtained by solving relaxations Boyd and Vandenberghe (2004); Geoffrion (2009) or by branch-and-bound algorithms Wolsey and Nemhauser (1999). However, in real-life applications, these dual bounds may be too loose to be informative, thus leaving practitioners to evaluate solution quality by relying on their domain knowledge and experience, which is prone to errors and may lead to sub-optimal decisions.
To address this challenge, this paper proposes a data-driven approach that evaluates the sub-optimality of a solution by leveraging uncertainty quantification (UQ) and conformal prediction (CP) techniques. Namely, the paper proposes to provide probabilistic guarantees for solution quality in the form of prediction intervals. To the best of the authors’ knowledge, this paper is the first to propose the use of CP to estimate solution sub-optimality in an optimization setting. In doing so, the paper introduces CPUL, a general framework for CP with upper and lower bounds. This contrasts with existing approaches from the optimization literature, which almost exclusively rely on computing valid, but possibly loose, dual bounds. The paper’s setting also contrasts with traditional CP settings that rely on approximate predictors, e.g., mean predictors Vovk et al. (2005) or quantile regression models Romano et al. (2019) which, in the present setting, may not yield valid lower/upper bounds.
@@ 1.1 Problem statement
where $x\in\mathcal{X}\subseteq\mathbb{R}^{p}$ denotes the instance’s parameters, $s\in\mathbb{R}^{n}$ are decision variables, $f_{x}:\mathbb{R}^{m}\rightarrow\mathbb{R}$ is the objective function to be minimized, $g_{x}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ encodes constraints and $\mathcal{S}_{x}=\{g_{x}(s)\leq 0\}$ is the set of feasible solutions, and $\Phi(x)$ is the instance’s optimal value. The paper assumes the availability of a primal-dual pair in the following form. A primal solution $\bar{s}\in\mathcal{S}_{x}$ provides a primal (upper) bound on the optimal value, i.e., $f_{x}(\bar{s})\geq\Phi(x)$. Primal solutions are obtained using either heuristics of exact algorithms. Conversely, a dual (lower) bound $\psi\in\mathbb{R}$ is guaranteed to be smaller than the optimal value, i.e., $\psi\leq\Phi(x)$. Dual bounds are obtained from relaxations and/or branch-and-bound algorithms, and provide a mathematical certificate on the quality of a primal solution $\bar{s}$. It is important to note that all modern global optimization algorithms return such a primal-dual pair. It is also possible to obtain a primal-dual pair using only machine learning (ML) models, by combining primal-feasible ML architectures Donti et al. (2021); Chen et al. (2023) with the dual lagrangian learning framework introduced in Tanneau and Hentenryck (2024).
@@
These predictors inherently form a 100% prediction interval $\hat{B}^{l}(X),\hat{B}^{u}(X)]$ for $Y$. Recall that, in a constrained optimization context, upper and lower predictors are obtained from primal and dual information, respectively. The goal of this paper is to derive tighter prediction intervals by targeting a $(1-\alpha)$ marginal coverage for $Y$ (where $\alpha\,{\in}\,[0,1]$), rather than perfect coverage.
Formally, consider a dataset of $N$ independent and identically distributed (i.i.d.) observations $(X_{i}\,{,}\,Y_{i})_{i\in[N]}$, where $X_{i}$ is sampled from an underlying distribution $\mathcal{P}$, and $Y_{i}\,{=}\,f(X_{i}),\forall i\,{\in}[N]$. Following the standard procedure of the split CP framework Vovk et al. (2005), divide the $N$ samples into a training and a calibration set, denoted as $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{cal}},$ with corresponding indices $\mathcal{I}_{\text{train}},\mathcal{I}_{\text{cal}}$. Next, consider a new i.i.d. observation $X_{N+1}$, sampled from $\mathcal{P}$. For $\alpha\,{\in}\,[0,1]$, the goal is to refine the initial prediction interval $\tilde{C}(X_{N+1})=[\hat{B}^{l}(X_{N+1}),\hat{B}^{u}(X_{N+1})]$ into a (smaller) prediction interval $\hat{C}(X_{N+1})=[\hat{L}(X_{N+1}),\hat{U}(X_{N+1})]$ that satisfies the marginal coverage guarantee
@@ 2 Background
UQ is crucial for informed decision making, particularly in real-world applications where prediction sets are preferred over point estimates. CP, first proposed by Vovk et al. (2005), is a widely used distribution-free UQ method, valued for its finite-sample coverage guarantees and computational efficiency. This section presents standard CP techniques, including Split CP, Conformal Quantile Regression, and Nested CP. While these methods can be applied to the problem at hand (see Section 5 and Appendix LABEL:app:experiments:cp_baselines), unlike the proposed framework, they are not designed to explicitly leverage valid lower and upper bound predictors.
@@
Vovk et al. (2005); Papadopoulos et al. (2002); Lei et al. (2018) is one of the most commonly used CP frameworks. Therein, a mean prediction model $\hat{f}$ is first trained using the training data $\mathcal{D}_{train}$. SCP then produces prediction intervals of the form
@@
where $\hat{s}\,{:}\,\mathbb{R}^{d}{\times}\mathbb{R}\,{\rightarrow}\,\mathbb{R}$ is a so-called conformity score, and $\hat{Q}^{\text{SCP}}_{\frac{\alpha}{2}},\hat{Q}^{\text{SCP}}_{1-\frac{\alpha}{2}}$, denote the empirical $\frac{\alpha}{2}$ and $1{-}\frac{\alpha}{2}$ quantiles of the scores $\{\hat{s}(X_{i},Y_{i})\}_{i\in\mathcal{I}_{\text{cal}}}$. Under the exchangeability assumption, the resulting prediction intervals have valid marginal coverage Vovk et al. (2005), i.e.,
@@
Common choices of conformity scores include the residual score $\hat{s}(x,y)\,{=}\,y{-}\hat{f}(x)$ and the absolute residual score $\hat{s}(x,y)\,{=}\,|y{-}\hat{f}(x)|$; the reader is referred to Angelopoulos and Bates (2021); Oliveira et al. (2024) for a more exhaustive review of SCP and conformity scores. The SCP methodology can be applied to the paper’s setting by replacing $\hat{f}$ in the above construction with either the lower ($\hat{B}^{l}$) or upper bound ($\hat{B}^{u}$) predictor. However, this fails to jointly utilize information from both bounds.
@@
To alleviate SCP’s lack of local adaptivity, CQR Romano et al. (2019) combines quantile regression models with a conformalization procedure: First, CQR trains quantile regression models to predict the quantiles $\alpha_{\text{lo}}$ and $\alpha_{\text{hi}}$. Then, conformalize as follows
@@
Gupta et al. (2022) introduces a unifying CP framework by nested prediction sets. NCP considers a family of nested prediction intervals $\{\hat{C}_{t}\}_{t\in\mathcal{T}\subseteq\mathbb{R}}$, i.e.,
@@
is computed from the calibration set. This framework encompasses all types of conformity scores, while maintaining the theoretical guarantees of standard CP Gupta et al. (2022). For instance, the CQR construction can be cast in the NCP framework by considering
@@
CP techniques have been use to analyze the precision of sketching algorithms, which are increasingly vital tools for handling massive datasets in modern machine learning Broder and Mitzenmacher (2004); Goyal et al. (2012); Cormode et al. (2018); Zhang et al. (2014). Sesia et al. (2023) construct prediction intervals for the frequency of a queried object based on a sketch Charikar et al. (2002). Therein, a deterministic upper bound $\hat{B}^{u}$ provided by Cormode and Muthukrishnan (2005) is combined with a trivial lower bound $\hat{B}^{l}=0$ (reflecting the non-negativity of counts) to construct nested intervals. The construction, referred to as “SFD CP” in this paper, is defined as:
@@
Sesia et al. (2023) also introduces an adaptive version of this construction, designed to better account for heteroscedasticity in residuals. The CPUL framework generalizes this approach by combining information from $\hat{B}^{l},\hat{B}^{u}$ in multiple ways; see Section 3.2.
Another related line of work leverages information from multiple models fitted on the training set. For example, Liang et al. (2024); Yang and Kuchibhotla (2024) focus on selecting the model that produces the most efficient prediction intervals. Extending these ideas, the CPUL framework reinterprets the problem in the context of multiple models, simultaneously utilizing both upper and lower bounds to maximize the utility of the available information.
@@ 3 CP With Upper And Lower Bounds
(b) Illustration of Paradoxical Miscoverage (motivation for OMLT): using a constant offset ($\pm t$) in the NCP construction results in an empty prediction interval where $\hat{B}^{u}(x)-\hat{B}^{l}(x)$ is small.
@@ 3.1 Exploiting valid bounds in NCP
Recall that $\hat{B}^{l},\hat{B}^{u}$ provide valid lower and upper bounds on the target variable, i.e., $\hat{B}^{l}(X)\,{\leq}\,Y\,{\leq}\,\hat{B}^{u}(X)$ always holds. This suggests a simple procedure for strengthening $\hat{C}$ without loss of coverage (as shown by Proposition 1) into
@@
$\hat{C}(x)\,{\cap}\,\big[\hat{B}^{l}(x),\hat{B}^{u}(x)\big],\forall x\in\mathcal{X}$. Then, $\mathbb{P}\big(Y_{N+1}\,{\in}\,\tilde{C}(X_{N+1})\big)=\mathbb{P}\big(Y_{N+1}\,{\in}\,\hat{C}(X_{N+1})\big)=1{-}\alpha.$
It is important to note that Proposition 1 holds irrespective of how the original prediction interval $\hat{C}$ is obtained. Theorem 1 shows that, if the calibration step is performed using the NCP framework, then there is no loss of performance whether the strengthening is performed before or after calibration.
Theorem 1.
Consider a family of nested prediction sets $\{\hat{C}_{t}\}_{t\in\mathcal{T}}$, where $\mathcal{T}\,{\subseteq}\,\mathbb{R}$, and let $\hat{\tau}$ be obtained following the NCP calibration step as per Eq. (11). Next, define the family of nested strengthened intervals $\{\tilde{C}_{t}\}_{t\in\mathcal{T}}$, where $\forall x\in\mathbb{R}^{d},\forall t\in\mathcal{T},\tilde{C}_{t}(x)=\hat{C}_{t}(x)\cap\big[\hat{B}^{l}(x),\hat{B}^{u}(x)\big],$ and let $\tilde{\tau}$ be obtained from Eq. (11). Then, $\hat{\tau}=\tilde{\tau}$ and $\forall x\in\mathbb{R}^{d},$
@@
Theorem 1 allows to decouple the calibration step from the strengthening. Thereby, one can perform the calibration step without access to $\hat{B}^{l}$ or $\hat{B}^{u}$, without any loss of coverage. This is particularly valuable when $\hat{B}^{l}$ or $\hat{B}^{u}$ are not available during training/calibration due to, e.g., privacy concerns.